. Consider Interest centers on the life of an electronic component. Suppose it is known that the probability that the component survives for more than 6000 hours is 0.42. Suppose also that the probability that the component survives no longer than 4000 hours is 0.04. Determine the Probability that;

(a) The life of the component is less than or equal to 6000 hours?

(b)The life is greater than 4000 hours? 


1
Expert's answer
2022-05-31T12:42:47-0400

(a) "P(X \le 6000)=1-P(X>6000)=1-0.42=0.58."

(b) "P(X>4000)=1-P(X\le4000)=1-0.04=0.96."